# A coupled-trajectory approach for decoherence, frustrated hops and internal consistency in surface hopping

### v1

@@ Abstract
A novel approach to coupling trajectories in surface hopping is presented. The coupled trajectory surface hopping algorithm based on the exact factorization of the molecular wavefunction is modified to allow energy sharing between the trajectories to avoid frustrated hops coupled with a deterministic hopping algorithm that gives more importance to the quantum populations. Promising results on a high dimensional model system suggest that this robust algorithm overcomes the common problems of overcoherence, frustrated hops and internal consistency in surface hopping.
@@
Nonadiabatic molecular dynamics is an irreplaceable method in the toolkit of physical chemists to model time-resolved photochemistry, the study of the coupled electron-nuclear evolution of molecules upon light excitation[1]. Among the many theories available nowadays, surface hopping, since its first appearance in the literature in 1971[2], has grown to become the method of choice to simulate the nonadiabatic dynamics of molecules and to accompany experimental observations.[3, 4, 5, 6, 7, 8, 9, 10, 11] Despite alternative theories being continuously developed, surface hopping remains very popular thanks to the unrivaled accuracy-cost compromise, to its relative implementation simplicity and especially to fact that it conveys a very intuitive picture of a nonadiabatic process.[12, 13, 14, 15, 16] In surface hopping, the nuclear wavepacket is represented in terms of a swarm of independent trajectories that classically follow Newton’s equations of motion on the support of electronic potential energy surfaces obtained as the solution of the time-independent Schrödinger equation at a given nuclear configuration. The key nonadiabatic effects, beyond the Born-Oppenheimer approximation, i.e., transitions between electronic states, are mimicked by stochastic hops of the trajectories from one potential energy surface to another usually occurring in localized regions of configuration space. Each trajectory of the swarm is coupled to the quantum-mechanical evolution of the electronic time-dependent state – even though not all surface hopping implementations require this.
@@
Surface hopping has its weaknesses, which have been reported extensively in the literature: it suffers from overcoherence, due to the disconnect between the dynamics of the classical nuclei and of the quantum electrons[17, 18, 19, 20, 21, 22, 16, 23, 24, 25]; when hops occur, energy conservation for the hopping trajectory is enforced, which causes frustrated hops if a trajectory hopping to a state of higher potential energy does not have enough kinetic energy to compensate for the increase in potential[26]; velocity rescaling related to the energy conservation after the hops is applied quite arbitrarily and strongly influence the dynamics;[27, 28, 29, 30, 31] using a swarm of independent trajectories to mimic the evolution of a quantum wavepacket misses non-local quantum-mechanical effects[19, 32, 33, 34, 35]. These weaknesses are consequences of the fact that surface hopping was not deduced originally from first principles, even though derivations [36, 32, 37] and qualitative justifications [38, 39] were proposed later on. Therefore, we ask ourselves, what is missing in surface hopping that would cure all these weaknesses at once? In this work, we address this question by using the idea of coupled trajectories emanating from the exact factorization [40, 41, 42, 43]. Note that a coupled-trajectory-based surface hopping scheme has been proposed previously [33, 34], but, as we will show below, it inherits some of the issues of the original surface hopping since the idea of coupled trajectories was not fully “embraced” by all aspects of the algorithm.
In this work, we propose a new strategy for a surface hopping procedure based on coupled trajectories and grounded on the exact factorization. Indeed, we are building upon previous work in this direction, that helped us understanding what is needed to propose a consistent, robust, accurate and efficient algorithm. Nonetheless, we believe that our proposition, which will be illustrated below, is an important step forward in the development of surface hopping schemes and is not to be regarded simply as a way to circumvent some of the well-known issues of surface hopping.
@@
It is worth discussing here the decoherence problem of surface hopping and its relation to the internal consistency. Internal consistency refers to the agreement between the two ways of estimating the population of the electronic states in surface hopping: the fraction of trajectories associated to, or evolving in, an electronic state (F); the trajectory-averaged population propagated along each trajectory (P), since usually a quantum-dynamical evolution equation for the electronic state is integrated along each trajectory. However, in general and unless decoherence corrections are used, it cannot be ensured that the two estimates agree with each other. The original hopping algorithm [2] ensures this internal consistency by using an average hopping probability over the swarm of trajectories for which the two estimates for the populations are the same. However, as soon as an independent trajectory formalism is invoked and the hopping probability is calculated at the level of a single trajectory, internal consistency is no longer guaranteed. The coupled-trajectory schemes that we propose here are naturally designed to reintroduce the idea of an average hopping probability to restore internal consistency. However, within CT-TSH this might introduce unphysical artifacts. Such an average hopping algorithm will assign the same hopping probability to all trajectories independently of their position. While this might ensure a correct distribution of the wavepacket/swarm with respect to the electronic states, it neglects the locality of the nonadiabatic couplings and the spatial distribution of the wavepacket, causing trajectories to undergo hops even in regions of negligible nonadiabatic coupling. This limitation has previously been overcome through the consensus surface hopping algorithm.[32] For CCT-TSH, we propose, by contrast, to forgo the stochastic hopping algorithm and introduce a deterministic hopping. [47] Since the electronic evolution is derived from the EF equation for the conditional amplitude, we drive the trajectory hops based on the values of the electronic populations predicted by this equation. In a way, the nuclear trajectories in CCT-TSH can be regarded merely as a support on which the time-dependent electronic state and the quantum momentum are evaluated. Therefore, it is crucial that they trajectories evolve following the electronic populations. For this reason, in CCT-TSH, the trajectories evolve in the electronic state with the largest population.
@@
Figure 2: Adiabatic population decays predicted by the different surface hopping schemes discussed in the main text. The lighter lines in each panel show the dynamics with different velocity rescaling procedures and treatments of frustrated hops, while the bold darker lines show the average of the individual runs. Solid lines show the average electronic populations (P) and dashed lines the fractions of trajectories (F).
@@
The number of frustrated hops during the dynamics varies strongly between the different methods (see Table 1). As expected, when we allow only the component of the velocity along the nonadiabatic coupling vectors to be rescaled, the number of frustrated hops is significantly large in all methods. Applying a decoherence correction as in TSH-ED, not only improves the internal consistency but also strongly decreases the number of frustrated hops. If the full velocity vector can be rescaled isotropically, there are no (or a negligible number of) frustrated hops during the dynamics. The number of frustrated hops during the CT-TSH dynamics is comparable to TSH, but surprisingly much larger if only the component of the velocity along the nonadiabatic coupling vectors can be rescaled. This issue is completely alleviated, albeit by construction, when using the equity-based and overlap-based CCT-TSH schemes. By contrast, the energy sharing scheme based on the quantum momentum shows a large number of frustrated hops. This can be explained by the modified hopping scheme where at each time step the energy sharing is attempted.
@@
In summary, we have identified a new and robust strategy to nonadiabatic dynamics using surface hopping. The key ingredient is the coupling of the trajectories, to achieve internal consistency, to circumvent overcoherence, to alleviate the frustrated hops, and to produce stable results upon changing the velocity rescaling. The coupling of the trajectories is essential to encode non-local, purely quantum mechanical features in the electronic and nuclear dynamics simulated with surface hopping, to calculate the quantum momentum and to introduce the energy sharing. Three practical schemes were proposed and tested on full-dimensional photodynamics of fulvene, based on a linear vibronic coupling model. Out of three, two schemes were shown to give robust results for electronic and nuclear properties. All the different flavors of CCT-TSH have been implemented with G-CTMQC [50], the LVC fulvene modes has been implemeted in QuantumModelLib[51], both are available upon request and will be made freely available on GitLab.

### v2

@@ Abstract
We address the issues of decoherence, frustrated hops and internal consistency in surface hopping. We demonstrate that moving away from an independent-trajectory picture is the strategy which allows us to propose a robust surface hopping scheme overcoming all these issues at once. Based on the exact factorization and on the idea of coupled trajectories, we consider the swarm of trajectories, that mimic the nuclear dynamics in nonadiabatic processes, as a unique entity. In this way, imposing energy conservation of the swarm and allowing the trajectories to share energy when hops occur clearly indicates the route towards a new surface hopping scheme. Encouraging results are reported, in terms of electronic and vibrational time-dependent properties on the photodynamics of fulvene, modeled with a full-dimensional linear vibronic coupling Hamiltonian.
@@
Nonadiabatic molecular dynamics is an irreplaceable method in the toolkit of physical chemists to model time-resolved photochemistry, the study of the coupled electron-nuclear evolution of molecules upon light excitation[1]. Among the many theories available nowadays, surface hopping, since its first appearance in the literature in 1971[2], has grown to become the method of choice to simulate the nonadiabatic dynamics of molecules and to accompany experimental observations.[3, 4, 5, 6, 7, 8, 9, 10, 11] Despite alternative theories being continuously developed, surface hopping remains very popular thanks to the unrivaled accuracy-cost compromise, to its relative implementation simplicity and especially to the fact that it conveys a very intuitive picture of a nonadiabatic process.[12, 13, 14, 15, 16, 17] In surface hopping, the nuclear wavepacket is represented in terms of a swarm of independent trajectories that classically evolve on the support of electronic potential energy surfaces obtained as the solution of the time-independent Schrödinger equation at a given nuclear configuration. The key nonadiabatic effects, beyond the Born-Oppenheimer approximation, i.e., transitions between electronic states, are mimicked by stochastic hops of the trajectories from one potential energy surface to another usually occurring in localized regions of configuration space. Each trajectory of the swarm is coupled to the quantum-mechanical evolution of the electronic time-dependent state – even though not all surface hopping implementations require this.[18, 19]
@@
Surface hopping has its weaknesses, which have been reported extensively in the literature: it suffers from overcoherence, due to the disconnect between the dynamics of the classical nuclei and of the quantum electrons[20, 21, 22, 23, 24, 25, 16, 26, 27, 28]; when hops occur, energy conservation for the hopping trajectory is enforced, which causes frustrated hops if a trajectory hopping to a state of higher potential energy does not have enough kinetic energy to compensate for the increase in potential[29]; velocity rescaling related to the energy conservation after the hops is applied quite arbitrarily and strongly influence the dynamics;[30, 31, 32, 33, 34] using a swarm of independent trajectories to mimic the evolution of a quantum wavepacket misses non-local quantum-mechanical effects[22, 35, 19, 36, 37]. These weaknesses are consequences of the fact that surface hopping was not deduced originally from first principles, even though derivations [38, 35, 39] and qualitative justifications [40, 41] were proposed later on.
In this work, we address the question: What is missing in surface hopping that would cure all its weaknesses at once? We demonstrate that the idea of coupled trajectories emanating from the exact factorization [42, 43, 44, 45] is capable to indicate the route to follow. Specifically, this route implies that the “surface hopping” idea is used as a tool to simplify conceptually and numerically the evolutions equations derived from the exact factorization, rather than as the working framework. In this way, we do not need to cure the weaknesses of surface hopping, since the starting point is a formulation which is naturally free from these weaknesses. Therefore, we propose a theoretical framework that fully exploits the coupling of the trajectories and that is combined with the algorithmic efficiency of surface hopping, bypassing introducing the issues of decoherence, frustrated hops and internal consistency. Note that a coupled-trajectory-based surface hopping scheme has been developed previously [19, 36], but, as we will show below, it inherits some of the issues of the original surface hopping since the idea of coupled trajectories was not fully “embraced” by all aspects of the algorithm. Instead, our new proposition is robust, stable and accurate, and, thus, we believe that our strategy is an important step forward in the development of novel ideas for nonadiabatic molecular dynamics schemes.
@@
It is worth discussing here the decoherence problem of surface hopping and its relation to the internal consistency. Internal consistency refers to the agreement between the two ways of estimating the population of the electronic states in surface hopping: the fraction of trajectories associated to, or evolving in, an electronic state (F); the trajectory-averaged population propagated along each trajectory (P), since usually a quantum-dynamical evolution equation for the electronic state is integrated along each trajectory. However, in general and unless decoherence corrections are used, it cannot be ensured that the two estimates agree with each other. The original hopping algorithm [2] ensures this internal consistency by using an average hopping probability over the swarm of trajectories for which the two estimates for the populations are the same. However, as soon as an independent trajectory formalism is invoked and the hopping probability is calculated at the level of a single trajectory, internal consistency is no longer guaranteed. The coupled-trajectory schemes that we propose here are naturally designed to reintroduce the idea of an average hopping probability to restore internal consistency. However, within CT-TSH this might introduce unphysical artifacts. Such an average hopping algorithm will assign the same hopping probability to all trajectories independently of their position. While this might ensure a correct distribution of the wavepacket/swarm with respect to the electronic states, it neglects the locality of the nonadiabatic couplings and the spatial distribution of the wavepacket, causing trajectories to undergo hops even in regions of negligible nonadiabatic coupling. This limitation has previously been overcome through the consensus surface hopping algorithm.[35] For CCT-TSH, we propose, by contrast, to forgo the stochastic hopping algorithm and introduce a deterministic hopping. [49] Since the electronic evolution is derived from the EF equation for the conditional amplitude, we drive the trajectory hops based on the values of the electronic populations predicted by this equation. In a way, the nuclear trajectories in CCT-TSH can be regarded merely as a support on which the time-dependent electronic state and the quantum momentum are evaluated. Therefore, it is crucial that the trajectories evolve following the electronic populations. For this reason, in CCT-TSH, the trajectories evolve in the electronic state with the largest population.
@@
Figure 2: Adiabatic population decays of S1 predicted by the different surface hopping schemes discussed in the main text. The lighter lines in each panel show the dynamics with different velocity rescaling procedures and treatments of frustrated hops, while the bold darker lines show the average of the individual runs. Solid lines show the average electronic populations (P) and dashed lines the fractions of trajectories (F).
@@
The number of frustrated hops during the dynamics varies strongly between the different methods (see Table 1). As expected, when we allow only the component of the velocity along the nonadiabatic coupling vectors to be rescaled, the number of frustrated hops is quite large in all methods. Applying a decoherence correction as in TSH-ED, not only improves the internal consistency but also strongly decreases the number of frustrated hops. If the full velocity vector can be rescaled isotropically, there are no (or a negligible number of) frustrated hops during the dynamics. The number of frustrated hops during the CT-TSH dynamics is comparable to TSH, but surprisingly much larger if only the component of the velocity along the nonadiabatic coupling vectors can be rescaled. This issue is completely alleviated, albeit by construction, when using the equity-based and overlap-based CCT-TSH schemes. By contrast, the energy sharing scheme based on the quantum momentum shows a large number of frustrated hops. This can be explained by the modified hopping scheme where at each time step the energy sharing is attempted.
@@
In summary, we have identified a new and robust strategy to simulate the electron-nuclear dynamics underlying nonadiabatic processes, which is based on the exact factorization and employs the idea of surface hopping to approximate the nuclear dynamics. The crucial ingredient is the coupling of the trajectories, a concept that emerges naturally in the context of the exact factorization, and is the key to achieve internal consistency, to circumvent overcoherence, to alleviate the frustrated hops, and to produce stable results upon changing the velocity rescaling within surface hopping. The coupling of the trajectories is essential to encode non-local, purely quantum mechanical features in the electronic and nuclear dynamics simulated using a surface hopping scheme, to calculate the quantum momentum and to introduce the energy sharing. Three practical schemes were proposed and tested on full-dimensional photodynamics of fulvene, based on a linear vibronic coupling model. Out of three, two schemes were shown to give robust results for electronic and nuclear properties. All the different flavors of CCT-TSH have been implemented in G-CTMQC [53] and are freely available on GitLab, while the LVC fulvene model has been implemented in QuantumModelLib[54].